Laval produces lamps and home lighting fixtures. Its most popular product is a brushed aluminum desk lamp. This lamp is made from components shaped in the fabricating department and assembled in the assembly department. Information related to the 28,000 desk lamps produced annually follows.

Direct materials $ 270,000
Direct labor
Fabricating department (6,500 DLH × $25 per DLH) $ 162,500
Assembly department (15,800 DLH × $26 per DLH) $ 410,800
Machine hours
Fabricating department 14,800 MH
Assembly department 20,450 MH

  
Expected overhead cost and related data for the two production departments follow.

Fabricating Assembly
Direct labor hours 140,000 DLH 285,000 DLH
Machine hours 173,000 MH 123,000 MH
Overhead cost $ 380,000 $ 430,000

  
Required
1. Determine the plantwide overhead rate for Laval using direct labor hours as a base.
2. Determine the total manufacturing cost per unit for the aluminum desk lamp using the plantwide overhead rate.
3. Compute departmental overhead rates based on machine hours in the fabricating department and direct labor hours in the assembly department.
4. Use departmental overhead rates from requirement 3 to determine the total manufacturing cost per unit for the aluminum desk lamps.
Answer #1

Solution f. Total direct Labor hours for both departments = 140000 + 285000 = 425000 Total overhead Cost = 380000 + 430000 =2) 270000 Computation of marfacturing Cost per unit particulars Amount Total manufacturing cost: Direct Material Derect Labor270000 162500 410800 Total Manufacturing cost : Direct Material Direct Labor: Fabricating blepacetiment Assembly department A